# The Impact of Visual Information in Chinese Characters: Evaluating Large Models’ Ability to Recognize and Utilize Radicals

### v1

@@ Abstract
The glyphic writing system of Chinese incorporates information-rich visual features in each character, such as radicals that provide hints about meaning or pronunciation. However, there has been no investigation into whether contemporary Large Language Models (LLMs) and Vision-Language Models (VLMs) can harness these sub-character features in Chinese through prompting. In this study, we establish a benchmark [^1] Our benchmark can be accessed through GitHub. to evaluate LLMs’ and VLMs’ understanding of visual elements in Chinese characters, including radicals, composition structures, strokes, and stroke counts. Our results reveal that models surprisingly exhibit some, but still limited, knowledge of the visual information, regardless of whether images of characters are provided. To incite models’ ability to use radicals, we further experiment with incorporating radicals into the prompts for Chinese language understanding tasks. We observe consistent improvement in Part-Of-Speech tagging when providing additional information about radicals, suggesting the potential to enhance CLP by integrating sub-character information.
@@ 4.2 Experimental Setup
We assess the visual information of Chinese characters using multilingual, bilingual, and open-source LLMs and VLMs. Multilingual LLMs include Aya-command Üstün et al. (2024), Claude-3 Anthropic (2024), Gemini-1.5, GPT-3.5 Turbo OpenAI , and GPT-4 OpenAI (2023). Chinese-English bilingual LLMs include ERNIE-Lite Baidu (2024a), Kimi-v1 MoonshotAI (2024), and open-source LLMs such as Baichuan-13B BaichuanInc (2024), BLOOM-7B BigScience (2024), ChatGLM-6B Zeng et al. (2023), Chinese-LLaMA-7B HFL (2024), InternLM-7B InternLM (2024), Orion-14B Chen et al. (2024), Qwen-7B Bai et al. (2023), Qwen-2-72B, and Yi-6B 01.AI et al. (2024). We also evaluate VLMs by providing images of query characters in the Microsoft YaHei [^9] YaHei is the default Chinese font in Microsoft Office. font to vision-capable models, including Claude-3, Gemini-1.5, GPT-4, and bilingual models Ernie-4 Baidu (2024b) and Kimi-v1. Additionally, we assess the pixel-based encoder model, PIXEL Rust et al. (2023). Since PIXEL is limited to specific tasks such as span-based QA, it is evaluated only on the multiple-choice structure recognition task after fine-tuning. To explore the models’ ability to learn Chinese visual information, we apply Chain-of-Thought (CoT) prompting and fine-tuning to GPT-3.5, and few-shot settings to both GPT-3.5 and GPT-4. The remaining models are evaluated in a zero-shot setting. We evaluate each result five times to calculate entropy for each task as an indicator of confidence. Detailed prompting and fine-tuning procedures are provided in Appendix C.1.
@@ 5.2 Experimental Setup
We select a series of LLMs for evaluation, including Aya-command, Claude-3, ERNIE-Lite-8K, GPT-3.5, GPT-4, and QWen-1.5 72B Chat. The models are instructed to return answers in JSON format, with target sentences annotated in a manner similar to Blevins et al. (2023). Each task and dataset is evaluated using 2,000 sample sentences five times. Due to higher costs, Claude-3 and GPT-4 are evaluated with 1,000 samples. We experiment with three different prompting methods:
@@ C.1 Detail Settings
For our evaluation, we use different sampling methods and settings based on the type of model. For LLMs, a random sample of 1,000 characters is selected for each task and model. Due to higher costs, the number of samples for VLMs is reduced to 500. ERNIE-V and Kimi-V, which lack API access, are tested manually with only 100 samples. We incorporate few-shot learning by providing models with three examples for each task, except for the structure recognition task, where one example per structure type is given. In the Chain-of-Thought (CoT) setting, models are prompted to break down their reasoning process step-by-step, with detailed prompts provided in the Appendix C.2. Models with fine-tuning are trained with a 7:3 split and tested using 1,000 samples randomly selected from the test set. To assess consistency and model entropy, each question is asked five times, and the best trial out of the five for each task is selected to calculate the overall results.
@@ C.4 GPT-4o and GPT-o1 Behavior
When simply prompting GPT-4o and GPT-4.1 with the question "What does this character mean?" for unfamiliar characters, both models provide responses with radical information, but they do so incorrectly (as shown in Figures 22 and 23). This suggests that while the models have been exposed to radicals in their training corpus, they are still unable to correctly identify or interpret the appropriate radicals.
@@ F.5 Aya Model Behavior
Figure 22: An example of GPT-o1’s response, with incorrect statements highlighted in red.
Figure 23: An example of GPT-4o’s response, with incorrect statements highlighted in red.

### v2

@@ Abstract
The glyphic writing system of Chinese incorporates information-rich visual features in each character, such as radicals that provide hints about meaning or pronunciation. However, there has been no investigation into whether contemporary Large Language Models (LLMs) and Vision-Language Models (VLMs) can harness these sub-character features in Chinese through prompting. In this study, we establish a benchmark [^1] Our benchmark can be accessed through GitHub. to evaluate LLMs’ and VLMs’ understanding of visual elements in Chinese characters, including radicals, composition structures, strokes, and stroke counts. Our results reveal that models surprisingly exhibit some, but still limited, knowledge of the visual information, regardless of whether images of characters are provided. To incite models’ ability to use radicals, we further experiment with incorporating radicals into the prompts for Chinese language processing (CLP) tasks. We observe consistent improvement in Part-Of-Speech tagging when providing additional information about radicals, suggesting the potential to enhance CLP by integrating sub-character information.
@@ 4.2 Experimental Setup
We assess the visual information of Chinese characters using multilingual, bilingual, and open-source LLMs and VLMs. Multilingual LLMs include Aya-command Üstün et al. (2024), Claude-3 Anthropic (2024), Gemini-1.5, GPT-3.5 Turbo OpenAI (2024a), and GPT-4 OpenAI (2023). Chinese-English bilingual LLMs include ERNIE-Lite Baidu (2024a), Kimi-v1 MoonshotAI (2024), and open-source LLMs such as Baichuan-13B BaichuanInc (2024), BLOOM-7B BigScience (2024), ChatGLM-6B Zeng et al. (2023), Chinese-LLaMA-7B HFL (2024), InternLM-7B InternLM (2024), Orion-14B Chen et al. (2024), Qwen-7B Bai et al. (2023), Qwen-2-72B, and Yi-6B 01.AI et al. (2024). We also evaluate VLMs by providing images of query characters in the Microsoft YaHei [^9] YaHei is the default Chinese font in Microsoft Office. font to vision-capable models, including Claude-3, Gemini-1.5, GPT-4, GPT-4o OpenAI (2024b), and bilingual models Ernie-4 Baidu (2024b) and Kimi-v1. Additionally, we assess the pixel-based encoder model, PIXEL Rust et al. (2023). Since PIXEL is limited to specific tasks such as span-based QA, it is evaluated only on the multiple-choice structure recognition task after fine-tuning. To explore the models’ ability to learn Chinese visual information, we apply Chain-of-Thought (CoT) prompting and fine-tuning to GPT-3.5, and few-shot settings to both GPT-3.5 and GPT-4. The remaining models are evaluated in a zero-shot setting. We repeat the evaluation for each task five times to compute entropy, using it as an indicator of the models’ confidence. Detailed prompting and fine-tuning procedures are provided in Appendix C.1.
@@ 5.2 Experimental Setup
We select a series of LLMs for evaluation, including Aya-command, Claude-3, ERNIE-Lite-8K, GPT-3.5, GPT-4, and QWen-1.5 72B Chat. The models are instructed to return answers in JSON format, with target sentences annotated in a manner similar to Blevins et al. (2023). Each task and dataset is evaluated using 2,000 sample sentences five times. Due to higher costs, Claude-3 and GPT-4 are evaluated with 1,000 samples. Additionally, we experiment with o1-mini OpenAI (2024c) using 100 samples, repeated three times. We experiment with three different prompting methods:
@@ C.1 Detail Settings
For our evaluation, we use different sampling methods and settings based on the type of model. For LLMs, a random sample of 1,000 characters is selected for each task and model. Due to higher costs, the number of samples for VLMs is reduced to 500. ERNIE-V and Kimi-V, which lack API access, are tested manually with only 100 samples. We incorporate few-shot learning by providing models with three examples for each task, except for the structure recognition task, where one example per structure type is given. In the Chain-of-Thought (CoT) setting, models are prompted to break down their reasoning process step-by-step, with detailed prompts provided in the Appendix C.2. Models with fine-tuning are trained in OpenAI’s platform using cross-entropy loss with a 7:3 split and tested using 1,000 samples randomly selected from the CCD dataset. To assess consistency and model entropy, each question is asked five times, and the best trial out of the five for each task is selected to calculate the overall results.
@@ C.4 GPT-4o and GPT-o1 Behavior
When simply prompting GPT-4o and GPT-4.1 with the question "What does this character mean?" for unfamiliar characters, both models provide responses with radical information, but they do so incorrectly (as shown in Figures 22). This suggests that while the models have been exposed to radicals in their training corpus, they are still unable to correctly identify or interpret the appropriate radicals.
@@ F.5 Aya Model Behavior
Figure 22: An example of GPT-4o and o1-preview’s response, with incorrect statements highlighted in red.